OpenStudy (anonymous):

is this a function yes or no? x=5y^(2)

OpenStudy (anonymous):

What is the definition of a function? Is there a test you know of to tell if something is a function or not?

OpenStudy (campbell_st):

nope... as it fails the vertical line test and each x value maps to 2 y values 

OpenStudy (anonymous):

You need to graph the question; by definition, the line so formed should not cross vertical line more than once. So, it's not.

OpenStudy (anonymous):

process of showing the relationship btwn the input and out of a problem.

OpenStudy (anonymous):

thank you all of you for all your help!

OpenStudy (goformit100):

Yo Yo Conic Sections.

OpenStudy (goformit100):

Thamks @Callisto

OpenStudy (anonymous):

You can also sometimes see this using algebra. The equation is given with x isolated, but normally standard form of equations has y, or f(x), isolated. If you isolate for y \[y ^{2} = \frac{ x }{ 5 }\]\[\pm y =\sqrt{\frac{ x }{ 5 }}\]\[y = \pm \sqrt{\frac{ x }{ 5 }}\]That plus/minus in from of the y, which happens whenever you "undo" a squared variable, means there are two options, a negative and a positive. If y is a negative number, then it becomes positive when it is squared.